OpenStudy (anonymous):

FAN/Medal - Factor completely: 2x^3 + 6x^2 + 10x + 30 2(x^3 + 3x^2 + 5x + 15) 2[(x^2 + 5)(x + 3)] (2x^2 + 10)(x + 3) (x^2 + 5)(2x + 6)

OpenStudy (anonymous):

i got x^2 +5 - 2x+3

OpenStudy (muzzack):

Reformatting the input : (1): "x2" was replaced by "x^2". 1 more similar replacement(s). Step by step solution : Step 1 : Simplify 2x3+6x2+10x + 30 Pulling out like terms : 1.1 Pull out like factors : 2x3 + 6x2 + 10x + 30 = 2 • (x3 + 3x2 + 5x + 15) Checking for a perfect cube : 1.2 x3 + 3x2 + 5x + 15 is not a perfect cube Trying to factor by pulling out : 1.3 Factoring: x3 + 3x2 + 5x + 15 Thoughtfully split the expression at hand into groups, each group having two terms : Group 1: 5x + 15 Group 2: x3 + 3x2 Pull out from each group separately : Group 1: (x + 3) • (5) Group 2: (x + 3) • (x2) ------------------- Add up the two groups : (x + 3) • (x2 + 5) Which is the desired factorization Polynomial Roots Calculator : 1.4 Find roots (zeroes) of F(x) = x2 + 5 Polynomial Roots Calculator is a set of methods aimed at finding values of x for which F(x)=0 Rational Roots Test is one of the above mentioned tools. It would only find Rational Roots that is numbers x which can be expressed as the quotient of two integers The Rational Root Theorem states that if a polynomial zeroes for a rational number P/Q then P is a factor of the Trailing Constant and Q is a factor of the Leading Coefficient In this case, the Leading Coefficient is 1 and the Trailing Constant is 5. The factor(s) are: of the Leading Coefficient : 1 of the Trailing Constant : 1 ,5 Let us test .... P Q P/Q F(P/Q) Divisor -1 1 -1.00 6.00 -5 1 -5.00 30.00 1 1 1.00 6.00 5 1 5.00 30.00 Polynomial Roots Calculator found no rational roots Final result : 2 • (x2 + 5) • (x + 3)

OpenStudy (muzzack):

excuse me \[2 \times (x ^{2} + 5) \times (x+3)\]

OpenStudy (anonymous):

so i get where we divide be 2 and get x3 + 3x2 + 5x + 15

OpenStudy (anonymous):

then split to 2 groups x3 + 3x2 and 5x + 15

OpenStudy (anonymous):

but got lost after that

OpenStudy (muzzack):

its B, just saying

OpenStudy (muzzack):

i replaced x^? to x? just to make it simple

OpenStudy (muzzack):

Thoughtfully split the expression at hand into groups, each group having two terms : Group 1: 5x + 15 Group 2: x3 + 3x2 Pull out from each group separately : Group 1: (x + 3) • (5) Group 2: (x + 3) • (x2) you have to make the y value equal to the other group

OpenStudy (anonymous):

ok ty

OpenStudy (muzzack):

yw :)

OpenStudy (anonymous):

Suppose \(f(x)=2x^{3} + 6x^{2} + 10x + 30\) Let \(f(x)=0\) So, A root of equation is \(x=-3\) => \(f(x)\) has a factor is \((x+3)\) dividing polynomial \(f(x)\) by \((x+3)\) we get \(2(x^{2}+5)\) Thus, \(2x^{3} + 6x^{2} + 10x + 30\) = \(2(x+3)\)\((x^{2}+5)\)